# SemGauss-SLAM: Dense Semantic Gaussian Splatting SLAM

### v2

@@ 3.3 Semantic-informed Bundle Adjustment
Currently, existing radiance field-based semantic SLAM systems utilize the latest input RGB-D frame to construct RGB and depth loss for pose estimation, while optimization of scene representation is then performed using estimated camera pose and the latest frame. However, relying solely on single-frame constraint for pose optimization can lead to cumulative drift in the tracking process due to the lack of global consistency. Furthermore, using only single-frame information for scene representation optimization can result in updates to the scene that are inconsistent on a global semantic level. To address this problem, we propose semantic-informed bundle adjustment, as illustrated in Fig. 3(b), to achieve joint optimization of 3D Gaussian representation and camera poses by leveraging multi-view constraints and semantic associations.
In semantic-informed bundle adjustment, we leverage the consistency of multi-view semantics to establish constraints. Specifically, rendered semantic feature is warped to its co-visible keyframe $j$ using estimated relative pose $T_{i}^{j}$, and constructs loss with extracted semantic feature $F_{j}^{e}$ of keyframe $j$ to obtain $\mathcal{L}_{\text{BA-sem}}$:
$\mathcal{L}_{\text{BA-sem}}=\sum_{i=1}^{N-1}\sum_{j=i+1}^{N}{(|T_{i}^{j}\cdot\mathcal{G}(T_{i},e)-F_{j}^{e}|)},$ (8)
where $\mathcal{G}(T_{i},e)$ represents splatted semantic embedding from 3D Gaussian $\mathcal{G}$ using camera pose $T_{i}$. Moreover, to achieve geometry and appearance consistency, we also warp rendered RGB and depth to co-visible keyframes to construct loss with corresponding keyframes following the similar approach in Eq.(8):
$\mathcal{L}_{\text{BA-rgb}}$ $=\sum_{i=1}^{N-1}\sum_{j=i+1}^{N}{(|T_{i}^{j}\cdot\mathcal{G}(T_{i},c)-F_{j}^{c}|)},$
$\mathcal{L}_{\text{BA-depth}}$ $=\sum_{i=1}^{N-1}\sum_{j=i+1}^{N}{(|T_{i}^{j}\cdot\mathcal{G}(T_{i},d)-F_{j}^{d}|)},$ (9)
where $F_{j}^{c}$ and $F_{j}^{d}$ are color and depth of keyframe $j$ respectively. $\mathcal{G}(T_{i},c)$ and $\mathcal{G}(T_{i},d)$ represents rendered RGB and depth. Therefore, overall loss function $\mathcal{L}_{\text{BA}}$ for joint optimization of corresponding keyframe poses and 3D Gaussian scene representation is the weighted sum of the above losses:
@@
where $\lambda_{e}$, $\lambda_{c}$, $\lambda_{d}$ are weighting coefficients. This design leverages fast rendering capability of 3D Gaussian to achieve joint optimization of scene representation and camera poses in real time. Furthermore, semantic-informed BA integrates consistency and correlation of multi-perspective semantic, geometry, and appearance information, resulting in low-drift tracking and consistent mapping.
Table 1: Quantitative comparison of tracking accuracy for our proposed SemGauss-SLAM with other dense visual SLAM and semantic SLAM methods on Replica dataset [21]. We utilize RMSE (cm) metric. To ensure more objectivity in the results, each scene is tested and averaged with three independent runs. Our work outperforms previous radiance field-based semantic SLAM methods across all scenes. Best results are highlighted as first, second.
Visual SLAM iMAP[23] 6.33 3.46 2.65 3.31 1.42 7.17 6.32 2.55 4.15
NICE-SLAM [33] 1.86 2.37 2.26 1.50 1.01 1.85 5.67 3.53 2.51
Vox-Fusion [29] 1.37 1.90 1.47 1.35 1.76 1.18 1.11 1.64 1.47
Co-SLAM [25] 0.72 0.85 1.02 0.69 0.56 2.12 1.62 0.87 1.06
ESLAM [6] 0.76 0.71 0.56 0.53 0.49 0.58 0.74 0.64 0.62
Point-SLAM [19] 0.61 0.41 0.37 0.38 0.48 0.54 0.69 0.72 0.52
SplaTAM [7] 0.31 0.40 0.29 0.47 0.27 0.29 0.32 0.55 0.36
Semantic SLAM SNI-SLAM [32] 0.50 0.55 0.45 0.35 0.41 0.33 0.62 0.50 0.46
DNS SLAM [9] 0.49 0.46 0.38 0.34 0.35 0.39 0.62 0.60 0.45
SemGauss-SLAM (Ours) 0.26 0.42 0.27 0.34 0.17 0.32 0.36 0.49 0.33
@@ 4.1 Experimental Setup
Datasets.   We evaluate the performance of SemGauss-SLAM on two datasets with semantic ground truth annotations, including 8 scenes on simulated dataset Replica [21] and 5 scenes on real-world dataset ScanNet [2].
Metrics.   To evaluate the SLAM system and rendering quality, we follow metrics from [19]. For mapping evaluation, we use Depth L1 (cm). For tracking accuracy evaluation, we utilize ATE RMSE (cm) [22]. Moreover, we use PSNR (dB), SSIM, and LPIPS for evaluating RGB image rendering performance. Semantic segmentation is evaluated with respect to mIoU accuracy.
Baselines.   We compare our method with the existing state of-the-art dense visual SLAM, including NeRF-based SLAM [23, 33, 25, 6, 19, 29] and 3D Gaussian SLAM [7]. For comparison of dense semantic SLAM performance, we consider NeRF-based semantic SLAM including SNI-SLAM [32], DNS SLAM [9], and NIDS-SLAM [4] as baseline.
Implementation Details.   We run SemGauss-SLAM on NVIDIA RTX 4090 GPU. For experimental settings, we perform mapping every 8 frames. The weighting coefficients of each loss are $\lambda_{f_{m}}=0.01$, $\lambda_{s_{m}}=0.01$, $\lambda_{c_{m}}=0.5$, $\lambda_{d_{m}}=1$ in mapping, $\lambda_{c_{t}}=0.5$ and $\lambda_{d_{t}}=1$ in tracking. Moreover, we set $\lambda_{e}=0.004$, $\lambda_{c}=0.5$, $\lambda_{c}=1$ in semantic-informed bundle adjustment. Please refer to the supplementary for further details of our implementation.
@@ 4.2 Experimental Results
Tracking Results.   As shown in Tab. 1 and 2, our method achieves up to 47% relative increase in tracking accuracy compared with other dense semantic SLAM methods. This improvement is attributed to semantic-informed bundle adjustment, which provides multi-view constraint leveraging consistency of semantic information to reduce accumulated drift in tracking process.
Table 2: We compare our proposed SemGauss-SLAM with other existing radiance field-based SLAM methods on ScanNet dataset [2] for tracking metric RMSE (cm). The results are an average of three independent runs. Best results are highlighted as first, second, third.
Methods scene0000 scene0059 scene0169 scene0181 scene0207 Avg.
NICE-SLAM [33] 12.00 14.00 10.90 13.40 6.20 11.30
Vox-Fusion [29] 68.84 24.18 27.28 23.30 9.41 30.60
Point-SLAM [19] 10.24 7.81 22.16 14.77 9.54 12.90
SplaTAM [7] 12.83 10.10 12.08 11.10 7.46 10.71
SemGauss-SLAM (Ours) 12.56 7.97 9.05 9.78 8.97 9.67
Rendering Quality Results.   Tab. 3 shows rendering quality on the input views of Replica dataset [21]. Our method achieves the best performance in PSNR, SSIM, and LPIPS compared with other dense semantic SLAM.
Table 3: Quantitative comparison of training view rendering performance on Replica dataset [21]. Our work outperforms other radiance field-based semantic SLAM methods on all three metrics across all scenes.
Methods Metrics room0 room1 room2 office0 office1 office2 office3 office4 Avg.
Visual SLAM NICE-SLAM [33] PSNR $\uparrow$ 22.12 22.47 24.52 29.07 30.34 19.66 22.23 24.94 24.42
SSIM $\uparrow$ 0.689 0.757 0.814 0.874 0.886 0.797 0.801 0.856 0.809
LPIPS $\downarrow$ 0.330 0.271 0.208 0.229 0.181 0.235 0.209 0.198 0.233
Vox-Fusion [29] PSNR $\uparrow$ 22.39 22.36 23.92 27.79 29.83 20.33 23.47 25.21 24.41
SSIM $\uparrow$ 0.683 0.751 0.798 0.857 0.876 0.794 0.803 0.847 0.801
LPIPS $\downarrow$ 0.303 0.269 0.234 0.241 0.184 0.243 0.213 0.199 0.236
Co-SLAM [25] PSNR $\uparrow$ 27.27 28.45 29.06 34.14 34.87 28.43 28.76 30.91 30.24
SSIM $\uparrow$ 0.910 0.909 0.932 0.961 0.969 0.938 0.941 0.955 0.939
LPIPS $\downarrow$ 0.324 0.294 0.266 0.209 0.196 0.258 0.229 0.236 0.252
ESLAM [6] PSNR $\uparrow$ 25.32 27.77 29.08 33.71 30.20 28.09 28.77 29.71 29.08
SSIM $\uparrow$ 0.875 0.902 0.932 0.960 0.923 0.943 0.948 0.945 0.929
LPIPS $\downarrow$ 0.313 0.298 0.248 0.184 0.228 0.241 0.196 0.204 0.239
SplaTAM [7] PSNR $\uparrow$ 32.86 33.89 35.25 38.26 39.17 31.97 29.70 31.81 34.11
SSIM $\uparrow$ 0.978 0.969 0.979 0.977 0.978 0.968 0.949 0.949 0.968
LPIPS $\downarrow$ 0.072 0.103 0.081 0.092 0.093 0.102 0.121 0.152 0.102
Semantic SLAM SNI-SLAM [32] PSNR $\uparrow$ 25.91 28.17 29.15 33.86 30.34 29.10 29.02 29.87 29.43
SSIM $\uparrow$ 0.885 0.910 0.938 0.965 0.927 0.950 0.950 0.952 0.935
LPIPS $\downarrow$ 0.307 0.292 0.245 0.182 0.225 0.238 0.192 0.198 0.235
SemGauss-SLAM (Ours) PSNR $\uparrow$ 32.55 33.92 35.15 39.18 39.87 32.97 31.60 35.00 35.03
SSIM $\uparrow$ 0.979 0.979 0.987 0.989 0.990 0.979 0.972 0.978 0.982
LPIPS $\downarrow$ 0.055 0.054 0.045 0.048 0.050 0.069 0.078 0.093 0.062
Reconstruction Results.   As shown in Tab. 4, our method achieves up to 77.3% increase in reconstruction accuracy compared with other radiance field-based methods. This enhancement in performance is due to the incorporation of semantic-informed BA process, achieving joint optimization for both camera poses and scene representation. Specifically, our method exploits the geometric consistency of co-visible frames to construct geometry constraints for more precise reconstruction.
Table 4: We compare our proposed SemGauss-SLAM with other existing radiance field-based SLAM methods on Replica [21] for reconstruction metric Depth L1 (cm). The results are an average of three independent runs. Our method outperforms other methods by up to 77.3%.
@@
Visual SLAM NICE-SLAM [33] 1.81 1.44 2.04 1.39 1.76 8.33 4.99 2.01 2.97
Vox-Fusion [29] 1.09 1.90 2.21 2.32 3.40 4.19 2.96 1.61 2.46
Co-SLAM [25] 1.05 0.85 2.37 1.24 1.48 1.86 1.66 1.54 1.51
ESLAM [6] 0.73 0.74 1.26 0.71 1.02 0.93 1.03 1.18 0.95
Semantic SLAM SNI-SLAM [32] 0.55 0.58 0.87 0.55 0.97 0.89 0.75 0.97 0.77
SemGauss-SLAM (Ours) 0.54 0.46 0.43 0.29 0.22 0.51 0.98 0.56 0.50
Semantic Segmentation Results.   As shown in Tab. 5, our work outperforms other dense semantic SLAM methods by up to 26% on mIoU metric and achieves 95% mIoU accuracy. Such enhancement attributes to the integration of semantic feature embedding into 3D Gaussian for enriched semantic representation, and semantic feature-level loss for direct guidance of semantic optimization. Moreover, our proposed semantic-informed BA also contributes to high semantic precision, as it leverages multiple co-visible frames to construct a globally consistent semantic map, thereby achieving high-precision semantic representation.
Table 5: Quantitative comparison of input views semantic segmentation performance on Replica [21] for semantic metric mIoU(%). For one scene, we calculate mIoU between splatted semantic maps and ground truth labels, which is calculated every 4 frames, to obtain average mIoU. Our method achieves highest semantic accuracy across all scenes.
Methods room0 room1 room2 office0 office1 office2 office3 office4
NIDS-SLAM [4] 82.45 84.08 76.99 85.94
DNS SLAM [9] 88.32 84.90 81.20 84.66
SNI-SLAM [32] 88.42 87.43 86.16 87.63 78.63 86.49 74.01 80.22
SemGauss-SLAM (Ours) 92.81 94.10 94.72 95.23 90.11 94.93 92.93 94.82
Table 6: Quantitative comparison of semantic novel view synthesis performance on Replica [21] for semantic metric mIoU(%). For one scene, we randomly choose 100 new viewpoints for semantic novel view synthesis evaluation. Our method outperforms SNI-SLAM [32] across all scenes, showing excellent 3D semantic mapping accuracy.
SNI-SLAM [32] 51.20 50.10 54.80 70.21 63.47 58.91 64.31 71.04 60.51
@@
Novel View Semantic Evaluation Results.   As shown in Tab. 6, our method achieves up to 75% increase in semantic segmentation accuracy for semantic novel view synthesis compared with existing dense semantic SLAM. By introducing 3D Gaussian feature embedding for semantic representation, our method enables continuous semantic modeling. This modeling is crucial for generating semantically coherent scenes, as it reduces the occurrence of sharp transitions that can cause inconsistencies, ensuring semantic representation accuracy from different viewpoints for high-precision novel view semantic synthesis.
Figure 4: Qualitative comparison on rendering quality of our method and baseline. We visualize 4 selected scenes of Replica dataset[21] and details are highlighted with red color boxes. Our method achieves photo-realistic rendering quality compared with other methods, especially in areas with rich textural information.
Visualization.   Fig. 4 shows rendering quality comparison of 4 scenes with interesting regions highlighted with colored boxes. For small objects, such as vases and items on the table, other methods fail to reconstruct them clearly. Our method leverages the high-quality rendering capability of 3D Gaussian representation and introduces semantic-informed BA to achieve detailed geometric reconstruction results. Specifically, our method enhances multi-view geometry consistency by BA to ensure that the reconstructed geometry aligns accurately across all observed viewpoints, leading to precise geometry reconstruction. As shown in Fig. 5, our method achieves superior novel view semantic segmentation accuracy compared with the baseline SNI-SLAM [32]. It can be observed that SNI-SLAM struggles with segmenting ceilings in novel view synthesis as they are less frequently observed during the mapping process. Consequently, the ceiling features are poorly modeled in the semantic scene representation. Our method introduces semantic-informed BA to construct multi-view constraints, enabling areas that are sparsely observed to effectively utilize the limited information from co-visible frames, thus establishing sufficient constraints for accurate semantic reconstruction.
Figure 5: Qualitative comparison on semantic novel view synthesis of our method and baseline semantic SLAM method SNI-SLAM [32] on 4 scenes of Replica[21].
Figure 6: Semantic rendering results and ground truth labels of feature-level loss ablation on two scenes of Replica [21].

### v3

@@ 3.3 Semantic-informed Bundle Adjustment
Currently, existing radiance field-based semantic SLAM systems utilize the latest input RGB-D frame to construct RGB and depth loss for pose estimation. Subsequently, the scene representation of these SLAM systems is optimized using the estimated pose and the latest frame. However, relying solely on single-frame constraint for pose optimization can lead to cumulative drift in the tracking process due to the absence of global constraints. Furthermore, using only single-frame information to optimize scene representation can result in globally inconsistent updates to the scene on the semantic level. To address this problem, we propose semantic-informed bundle adjustment to achieve joint optimization of 3D Gaussian representation and camera poses by leveraging multi-view constraints and semantic associations.
In semantic-informed BA, we leverage the consistency of multi-view semantics to establish constraints. Specifically, rendered semantic feature is warped to its co-visible frame $j$ using estimated relative pose transformation $T_{i}^{j}$, and constructs loss with rendered semantic feature $\mathcal{G}(T_{j},e)$ of frame $j$ to obtain $\mathcal{L}_{\text{BA-sem}}$:
$\mathcal{L}_{\text{BA-sem}}=\sum_{i=1}^{N-1}\sum_{j=i+1}^{N}{(|T_{i}^{j}\cdot\mathcal{G}(T_{i},e)-\mathcal{G}(T_{j},e)|)},$ (8)
where $\mathcal{G}(T_{i},e)$ represents splatted semantic embedding from 3D Gaussian $\mathcal{G}$ using camera pose $T_{i}$. Moreover, to achieve geometry and appearance consistency, we also warp rendered RGB and depth to co-visible frames to construct loss with corresponding frames following similar approach in Eq.(8):
$\mathcal{L}_{\text{BA-rgb}}$ $=\sum_{i=1}^{N-1}\sum_{j=i+1}^{N}{(|T_{i}^{j}\cdot\mathcal{G}(T_{i},c)-\mathcal{G}(T_{j},c)|)},$
$\mathcal{L}_{\text{BA-depth}}$ $=\sum_{i=1}^{N-1}\sum_{j=i+1}^{N}{(|T_{i}^{j}\cdot\mathcal{G}(T_{i},d)-\mathcal{G}(T_{j},d)|)},$ (9)
where $\mathcal{G}(T_{i},c)$ and $\mathcal{G}(T_{i},d)$ represents rendered RGB and depth of frame $i$. Therefore, overall loss function $\mathcal{L}_{\text{BA}}$ for joint optimization of corresponding frame poses and 3D Gaussian scene representation is the weighted sum of the above losses:
@@
where $\lambda_{e}$, $\lambda_{c}$, $\lambda_{d}$ are weighting coefficients. This design leverages fast rendering capability of 3D Gaussian to achieve joint optimization of scene representation and camera poses. Furthermore, semantic-informed BA integrates consistency and correlation of multi-perspective semantic, geometry, and appearance information, resulting in low-drift tracking and consistent mapping.
@@ 4.1 Experimental Setup
Datasets.   We evaluate the performance of SemGauss-SLAM on two datasets with semantic ground truth annotations, including 8 scenes on simulated dataset Replica straub2019replica and 5 scenes on real-world dataset ScanNet dai2017scannet .
Metrics.   We follow metrics from keetha2023splatam to evaluate the SLAM system accuracy and rendering quality. For reconstruction metric, we use Depth L1 (cm). For tracking accuracy evaluation, we use ATE RMSE (cm) sturm2012benchmark . Moreover, we use PSNR (dB), SSIM, and LPIPS for evaluating RGB image rendering performance. Semantic segmentation is evaluated with respect to mIoU (%) metric.
Baselines.   We compare our method with the existing state of-the-art dense visual SLAM, including NeRF-based SLAM zhu2022nice ; wang2023co ; eslam ; sandstrom2023point ; yang2022vox and 3DGS-based SLAM keetha2023splatam . For comparison of dense semantic SLAM performance, we consider NeRF-based semantic SLAM zhu2023sni ; li2023dns ; haghighi2023neural , and SGS-SLAM li2024sgs which is a work parallel to ours, as baseline.
Implementation Details.   We run SemGauss-SLAM on NVIDIA RTX 4090 GPU. For experimental settings, we perform mapping every 8 frames. The weighting coefficients of each loss are $\lambda_{f_{m}}=0.01$, $\lambda_{s_{m}}=0.01$, $\lambda_{c_{m}}=0.5$, $\lambda_{d_{m}}=1$ in mapping, $\lambda_{c_{t}}=0.5$ and $\lambda_{d_{t}}=1$ in tracking. Moreover, we set $\lambda_{e}=0.004$, $\lambda_{c}=0.5$, $\lambda_{c}=1$ in semantic-informed bundle adjustment.
Table 1: Quantitative comparison of SLAM accuracy and rendering performance for our proposed SemGauss-SLAM and other radiance field-based SLAM methods. The results are an average of 8 scenes on Replica dataset straub2019replica . To ensure the reliability of the results, each scene is tested and averaged with five independent runs. Our work outperforms previous works of all metrics.
Methods Reconstruction Localization Rendering Quality
Depth L1[cm] $\downarrow$ ATE RMSE[cm] $\downarrow$ PSNR(dB) $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
NICE-SLAM zhu2022nice 2.97 2.51 24.42 0.809 0.233
Vox-Fusion yang2022vox 2.46 1.47 24.41 0.801 0.236
Co-SLAM wang2023co 1.51 1.06 30.24 0.939 0.252
ESLAM eslam 0.95 0.62 29.08 0.929 0.239
SplaTAM keetha2023splatam 0.73 0.41 34.11 0.968 0.102
SNI-SLAM zhu2023sni 0.77 0.46 29.43 0.935 0.235
SemGauss-SLAM (Ours) 0.50 0.33 35.03 0.982 0.062
Figure 3: Qualitative comparison on rendering quality of our method and baseline. We visualize 4 selected scenes of Replica dataset straub2019replica and details are highlighted with red color boxes. Our method achieves photo-realistic rendering quality and higher completion of reconstruction, especially in areas with rich textural information.
@@ 4.2 Experimental Results
SLAM and Rendering Quality Results.   As shown in Tab. 1, our method achieves the highest accuracy in all metrics compared with other radiance field-based SLAM and up to 35% relative increase in reconstruction accuracy. This enhancement is attributed to the incorporation of semantic-informed bundle adjustment, which provides multi-view constraint to achieve joint optimization of
Table 2: We compare our proposed SemGauss-SLAM with other existing radiance field-based SLAM methods on ScanNet dataset dai2017scannet for tracking metric RMSE (cm) $\downarrow$.
Methods $0$ 0059 0169 0181 0207 Avg.
NICE-SLAM zhu2022nice 12.00 14.00 10.90 13.40 6.20 11.30
Vox-Fusion yang2022vox 68.84 24.18 27.28 23.30 9.41 30.60
Point-SLAM sandstrom2023point 10.24 7.81 22.16 14.77 9.54 12.90
SplaTAM keetha2023splatam 12.83 10.10 12.08 11.10 7.46 10.71
SemGauss-SLAM (Ours) 11.87 7.97 8.70 9.78 8.97 9.46
both camera poses and scene representation. Tab. 2 demonstrates our method outperforms baseline methods on real-world dataset ScanNet dai2017scannet .
Novel View Semantic Evaluation Results.   As shown in Tab. 3, our method achieves up to 49% relative increase in semantic novel view synthesis compared with existing dense semantic SLAM. By introducing 3D Gaussian feature embedding for semantic representation, our method enables continuous semantic modeling. This modeling is crucial for generating semantically coherent scenes, as it reduces the occurrence of sharp transitions that can cause inconsistencies, ensuring semantic representation accuracy from different viewpoints for high-precision novel view semantic synthesis.
Table 3: Quantitative comparison of semantic novel view synthesis performance on Replica straub2019replica for semantic metric mIoU(%). For one scene, we randomly choose 100 new viewpoints that are different from SLAM mapping perspectives for evaluation and mIoU is calculated between splatted semantic labels and ground truth labels. Our method outperforms baseline across all scenes, showing excellent 3D semantic mapping accuracy.
@@
SNI-SLAM zhu2023sni 51.20 50.10 54.80 70.21 63.47 58.91 64.31 71.04 60.51
@@
Table 4: Quantitative comparison of input views semantic segmentation performance on 4 scenes of Replica straub2019replica for semantic metric mIoU(%), as its results are only reported on such scenes. For one scene, we calculate mIoU between splatted semantic maps and ground truth labels every 4 frames to obtain average mIoU. Our method achieves the highest semantic accuracy across all scenes.
Methods room0 room1 room2 office0 Avg.
GT NIDS-SLAM haghighi2023neural 82.45 84.08 76.99 85.94 82.37
DNS SLAM li2023dns 88.32 84.90 81.20 84.66 84.77
SNI-SLAM zhu2023sni 88.42 87.43 86.16 87.63 87.41
SGS-SLAM li2024sgs 92.95 92.91 92.10 92.90 92.72
Ours (GT) 96.30 95.82 96.51 96.72 96.34
Seg Ours 92.81 94.10 94.72 95.23 94.22
Figure 4: Qualitative comparison on semantic novel view synthesis of our method and baseline semantic SLAM method SNI-SLAM zhu2023sni on 4 scenes of Replica straub2019replica .
Semantic Segmentation Results.   Since haghighi2023neural ; li2023dns ; li2024sgs ; zhu2023sni employs ground truth labels for supervision, for a fair comparison, we also present our results using ground truth labels for supervision, denoted as Our (GT). The results of using semantic segmentation results for supervision are Ours. As shown in Tab. 4, our work outperforms existing radiance field-based semantic SLAM methods. Such enhancement attributes to the integration of semantic feature embedding into 3D Gaussian for enriched semantic representation, and semantic feature-level loss for direct guidance of semantic optimization. Moreover, our proposed semantic-informed BA also contributes to high semantic precision, as it leverages multiple co-visible frames to construct a globally consistent semantic map for high-precision semantic representation.
Visualization.   Fig. 3 shows rendering quality comparison of 4 scenes with interesting regions highlighted with colored boxes. For less-frequent observed areas, such as floor and side of sofa, other methods either result in low-quality reconstruction or leave holes. Our method leverages the high-quality rendering capability of 3D Gaussian representation and introduces semantic-informed BA to achieve detailed and complete geometric reconstruction results. Specifically, our method enhances multi-view geometry consistency by BA to ensure that the reconstructed geometry aligns accurately across all observed viewpoints, leading to precise and complete geometry reconstruction. As shown in Fig. 4, our method achieves superior novel view semantic segmentation accuracy compared with baseline SNI-SLAM zhu2023sni . It can be observed that SNI-SLAM struggles with segmenting ceilings in novel view synthesis as they are less frequently observed during the mapping process. Consequently, the ceiling features are poorly modeled in the semantic scene representation. Our method introduces semantic-informed BA to construct multi-view constraints, enabling areas that are sparsely observed to effectively utilize the limited information from co-visible frames, thus establishing sufficient constraints for accurate semantic reconstruction.